HUBBLE SPACE TELESCOPE DAILY REPORT #3077 PERIOD COVERED: 0000Z (UTC) 03/20/02 - 0000Z (UTC) 03/21/02 Daily Status Report as of 080/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.2 Completed Four Sets of STIS/CCD 8963 (CCD Bias) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD read noise from multiple bias frames obtained in common science modes: amplifier D, gains 1 and 4 electrons/DN, and binning factors of 1x1, 1x2, 2x1, and 2x2. The read noise measurements before and after NCC turn-on are then compared, and standard calibration reference files for use in pipeline reduction of SMOV3B data are constructed. The proposal had no reported problems. 1.3 Completed STIS/CCD 8964 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain dark frames twice daily, as in a corresponding calibration program. Standard calibration reference darks are constructed and hot pixel lists are updated for use in pipeline reduction of SMOV3B data. All observations completed without incident. 1.4 Completed Two Sets of NCS 8967 (Start NICMOS Cooling System) The NICMOS Cooling System was configured to re-cool the NICMOS detectors, to determine the optimal science operating temperature for the NICMOS detectors, and to verify stable operation. The goals are as follows: 1} cool the NICMOS to the coldest possible temperature as specified by the NICMOS team {73K} with HST at a hot attitude and the NCC compressor operating near its maximum allowable speed. 2} measure NICMOS detector performance during the cool-down and steady state operation of the NCS. 3} demonstrate stability {+/-0.1K} of the NICMOS detector temperature at the optimal science operating temperature and 4} demonstrate repeatability {+/-0.1 K} of NICMOS detector temperature following changes from the optimal science operating temperature. All tests were successful. 1.5 Completed Nine Sets of NICMOS/1/2/3 8945 (Cooling System Monitoring) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3 was used 1}.to measure NICMOS detector performance during the cool-down and steady state operation of the NCS. 2}. to demonstrate stability {+/-0.1K} of the NICMOS detector temperature at the optimal science operating temperature. and 3}.to demonstrate repeatability {+/-0.1 K} of NICMOS detector temperature following changes from the optimal science operating temperature. No problems were encountered. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 0 Successful: 0 Scheduled Re-acquisitions: 0 Successful: 0 2.2 FHST Updates: Scheduled: 0 Successful: 0 2.3 Operations Notes: Real-time operations returned to the STScI SIMOR at 080/0230Z. The NICMOS Cryo Cooler continues to operate nominally. Several attitudes reference update slews were performed in order to maintain HST attitude. The following operations notes were utilized during the day: TPS limit change Limit violations due to increased load ACS limit changes Monitor ACS input voltage 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW